OpenStudy (anonymous):

Find the slope of the line that passes through each pair of points. (-1/4, 3/4), (1/2, -3/4) I'm not exactly sure what to do to be completely honest. Where do I start?

OpenStudy (anonymous):

@alyssa_xo do u know how to do this??

OpenStudy (alyssa_xo):

use the slope formula \(\huge m=\frac{\Delta y}{\Delta x}=\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (texaschic101):

slope formula : (y2 - y1) / (x2 - x1) (-1/4, 3/4) ---- x1 = -1/4 and y1 = 3/4 (1/2, -3/4) --- x2 = 1/2 and y2 = -3/4 Now sub slope = (-3/4 - 3/4) / (1/2 -(- 1/4) slope = (-6/4) / (1/2 + 1/4) slope = (-6/4) / (2/4 + 1/4) slope = (-3/2) / (3/4) slope = -3/2 * 4/3 slope = - 12/6 slope = -2

OpenStudy (anonymous):

@texaschic101 i'm a little confused with where the addition came in

OpenStudy (alyssa_xo):

subtraction of a negative = positive 1-(-3) = 4

OpenStudy (alyssa_xo):

you can view it as the negatives "cancelling"